OpenStudy (anonymous):

Evaluate the limit at -infinity. x/ (x-sqrt(x^2+7))

OpenStudy (anonymous):

i did it and yet couldnt find out

OpenStudy (anonymous):

\[\frac{x}{x -\sqrt{x^2- 7 }}\] \[ \frac{1}{1 - \sqrt{1 - \frac{7}{x^2}}}\]No...maybe some other approach..is the answer undefined?

OpenStudy (anonymous):

james i tried it already.. couldnt get the answer it should be 1/2

OpenStudy (turingtest):

I'm trying and I can't get it either!

OpenStudy (anonymous):

im so desperate .. its not homework or something i just want to do it!

OpenStudy (turingtest):

yes it comes out as undefined either way, right?

OpenStudy (turingtest):

or the book is wrong...

OpenStudy (anonymous):

i put it in "limit calculator" and it gives me 1/2 also when i plug really large numbers it gives me 1/2

OpenStudy (anonymous):

really large negative* numbers

myininaya (myininaya):

the limit is 1/2 i can show you

OpenStudy (anonymous):

ill be happy

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay, hmm if I think about the equation I do get \(\frac{1}{2}\)

OpenStudy (anonymous):

My TI-89 returns 1/2

OpenStudy (anonymous):

\[\sqrt{x^2+7} \to x \]Right?

myininaya (myininaya):

\[\lim_{x \rightarrow -\infty}\frac{x}{x-\sqrt{x^2+7}} \] since x<0=>|x|=-x remember \[|x|=\sqrt{x^2}\] so anyways top and bottom by \[\sqrt{x^2}\] \[\lim_{x \rightarrow -\infty} \frac{\frac{x}{\sqrt{x^2}}}{\frac{x}{\sqrt{x^2}}-\sqrt{\frac{x^2+7}{x^2}}}\] \[\lim_{x \rightarrow -\infty}\frac{\frac{x}{-x}}{\frac{x}{-x}-\sqrt{1+\frac{7}{x^2}}}\]

OpenStudy (anonymous):

And also keep in mind we have negative large numbers here ...

myininaya (myininaya):

\[\frac{-1}{-1-1}=\frac{-1}{-2}=\frac{1}{2}\]

OpenStudy (anonymous):

So it's kinds like \[\frac{-x}{-x - x}\to \frac{1}{2}\]

OpenStudy (anonymous):

very nice .. thank you all :)

OpenStudy (turingtest):

I used L'Hospital's Rule and got\[\lim_{x \rightarrow -\infty}1/[1-(1/\sqrt{1+7/x^2})\] which if you can figure an excuse to take the negative root in the denominator yields 1/2. That's just ad hoc though.

myininaya (myininaya):

you don't have to use l'hospital's rule but whatever floats your boat

OpenStudy (okagurao):

idkkkkkkkkkkkkkkkk